OpenStudy (anonymous):

let Y be a random variable with the mean , and standard deviation . Let W=2Y. compute the mean and standard deviation of W

OpenStudy (anonymous):

Need help?

OpenStudy (anonymous):

yyesss

OpenStudy (anonymous):

What do you have so far?

OpenStudy (anonymous):

um nothing. lol

OpenStudy (anonymous):

i just know various equations: sqare root of V(y)

OpenStudy (anonymous):

E(y)=Sumation (P(y))

OpenStudy (anonymous):

First, just curious: did they give you a specific value for E(Y) and \[\sigma _{Y}\]

OpenStudy (anonymous):

And, are you looking for an explanation or just the answer?

OpenStudy (anonymous):

no . i was only given the info that i told u

OpenStudy (anonymous):

preferably explanation with an answer

OpenStudy (anonymous):

It might be more helpful if I give you the answers first. E(W) = E(2Y) = 2*E(Y) Stdv(W) = Stdv(2Y) = 2*Stdv(Y)

OpenStudy (anonymous):

Think of it this way: Let's say you have a random group of men whose average height is 6ft. That's the same as saying their average height is 72 inches, right? Well, that's just 12*E(Y) = E( 12*Y) Does that part make sense? It's ok if it doesn't...

OpenStudy (anonymous):

Sometimes you see it written as: E(aX + b) = aE(X) + b where "X" is your random variable and "a" and "b" are constants.

OpenStudy (anonymous):

ya. it makes sense so far

OpenStudy (anonymous):

i was thinking E(w/2)=Y

OpenStudy (anonymous):

i really understand deviation and mean when using concrete examples. but what numbers do i plug in?

OpenStudy (anonymous):

Well, I think the point is to rewrite the expected value of W in terms of the expected value of Y. Also, don't just write "Y" -- you have to write "E(Y)" -- because Y is a random variable and, well, that's the kind of thing that teachers like to take points off for. Last thing: Remember, the standard deviation is ALWAYS POSITIVE. So if you have W = 2Y or W=-2Y, the standard deviation of W will be the same either way (i.e. it will equal 2*stdv(Y) ).

OpenStudy (anonymous):

Here's an example: Try calculating the the mean and standard deviation of a roll of a die. Then, instead of 1,2,3,...,6 multiply each number by 5 and then recalculate the mean and standard deviation. i.e. with the same probabilities but the numbers 5,10,15, ..., 30

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

5+10+15+20+25+30 / 6 ?

OpenStudy (anonymous):

Yeah, you could. Or, if you happen to know that the expected value of a regular die throw is 3.5, then you can just multiply 5*3.5 = 17.5 and skip all that extra addition and division.

OpenStudy (anonymous):

how do u know each is 3.5?

OpenStudy (anonymous):

(1+2+3+4+5+6)/6

OpenStudy (anonymous):

um okay? and that gives you the probability of each roll?

OpenStudy (anonymous):

isnt each roll just 1/6 probable?

OpenStudy (anonymous):

No, it gives the EXPECTED VALUE of each roll.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

okay so back to my example, how do i do it?

OpenStudy (anonymous):

The probability is the same for all the sides

OpenStudy (anonymous):

which example?

OpenStudy (anonymous):

lol my very abstract question

OpenStudy (anonymous):

the W=2Y?

OpenStudy (anonymous):

what do i do with 2y=w?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh yeah, ha! That one...

OpenStudy (anonymous):

Go back to the dice roll. That's like me saying: Roll a die. Take that number and multiply it by two. That's your random variable...

OpenStudy (anonymous):

so its 2y^2?

OpenStudy (anonymous):

E(2Y)=2y ^2 ?

OpenStudy (anonymous):

No, just 2*Y

OpenStudy (anonymous):

but you just said sqaure it?

OpenStudy (anonymous):

When? I thought I said "multiply it by two"

OpenStudy (anonymous):

ugh what do i do

OpenStudy (anonymous):

If you have a random variable "Y" then since W=2Y, that means that the value of W is two times the value of Y.

OpenStudy (anonymous):

okay so how do i find the mean ? and variance?